OpenStudy (psymon):

Reduction of Order: Professor didn't have time to go over it, but still assigned it for hw. Anyone mind working me through a problem?

OpenStudy (psymon):

\[x^{2}y'' - xy' + 2y = 0; y _{1}= xsin(lnx)\]

OpenStudy (psymon):

And sorry if I reply slow, working a bit. Any help is appreciated still.

OpenStudy (amistre64):

reduction of order tends to imply solving a u' equaivalent setup

OpenStudy (psymon):

Yeah, finding a way to makeit linear in u using a sub, correct?

OpenStudy (amistre64):

yep

OpenStudy (psymon):

I just dont know the steps in doing that, lol. My book is pathetic :/

OpenStudy (amistre64):

Pauls tends to be very readable to me http://tutorial.math.lamar.edu/Classes/DE/ReductionofOrder.aspx

OpenStudy (amistre64):

assume y2 = u y1

OpenStudy (psymon):

Right, so y2 = uxsin(lnx)

OpenStudy (amistre64):

and y2', y2'' from that gets us?

OpenStudy (psymon):

Oh joy. y' = u'xsin(lnx) + usin(lnx) + ucos(lnx) y'' = u''xsin(lnx) + u'sin(lnx) + u'cos(lnx) + u'sin(lnx) + ucos(lnx)/x + u'cos(lnx) - usin(lnx)/x I know Ill have to simplify that, assuming I did it right. But Yeah, plug it in... x^2(u''xsin(lnx)+2u'sin(lnx) + 2u'cos(lnx) + (u/x)[cos(lnx) - sin(lnx)) - x(u'xsin(lnx) + usin(lnx) + ucos(lnx)) + 2(uxsin(lnx) Yeah, thats awful.

OpenStudy (amistre64):

:) \[u''x^3\sin(\ln x)+2u'x^2[\sin(\ln x)+\cos(\ln x)]-ux[\sin(\ln x)-\cos(\ln x)]\\ ~~~~~~~~~~~~~~~~~~~~~~-u'x^2[\sin(\ln x)]~~~~~~~~~~~~~~~~~-ux[\sin(\ln x+\cos(\ln x))]\\ ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ +2ux[\sin(\ln x)] \]

OpenStudy (psymon):

Ill take your word for it. Dunno if you simplifie it more or if I just made a mistake in all that mess. Bit distracted, so forgive me xD

OpenStudy (amistre64):

\[u''x^3\sin(\ln x)+u'x^2\sin(\ln x)+2u'x^2\cos(\ln x)]+ux[\cos(\ln x)]\\ ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~-ux[\cos(\ln x))]\\\] \[u''x^3\sin(\ln x)+u'x^2[\sin(\ln x)+2\cos(\ln x)=0 \]

OpenStudy (psymon):

Nice :P

OpenStudy (amistre64):

since the term in u drops out, we can "reduce" this by an order: let w=u'

OpenStudy (amistre64):

looks like i forgot some brackets and what not in the latex code :) but i think its still readable

OpenStudy (amistre64):

looks like i forgot some brackets and what not in the latex code :) but i think its still readable

OpenStudy (amistre64):

\[w'x^3\sin(\ln x)+w~x^2[\sin(\ln x)+2\cos(\ln x)]=0\] \[w'\sin(\ln x)+\frac1xw[\sin(\ln x)+2\cos(\ln x)]=0\] \[w'+\frac1x[1+2cos(\ln x)]~w=0\]

OpenStudy (amistre64):

cotangent in that last one, not cos

OpenStudy (psymon):

The term in u drops out? Obviously not talking about derivatives. So is that just when the u term with no derivatives drops? And you said because of that, does that mean its supposed to happen or did we get lucky with something?

OpenStudy (amistre64):

its spose to happen in order to have the u'' and u' parts all alone

OpenStudy (amistre64):

since the name of a variable is not important, we can state this in terms of say: u'' = w'; and u' = w to "reduce" the order

OpenStudy (psymon):

So integrating factor is xsin^2(x)

OpenStudy (amistre64):

maybe ... with any luck :)

OpenStudy (amistre64):

cot, not cos

OpenStudy (psymon):

Yeah, Im pretty sure I did it with cot, lol.

OpenStudy (amistre64):

\[w'+\frac1x[1+2\cot(\ln x)]~w=0\]

OpenStudy (psymon):

So, assuming Im right, I would have just : \[\int\limits_{}^{}(wxsin^{2}x)'= 0 \] \[wxsin^{2}x = C\] Im just guessing that its supposed to do that, but im not 100% positive considering its 0 over there.

OpenStudy (psymon):

oops, meant the x in sin to be lnx, lol

OpenStudy (amistre64):

\[\int~\frac1x+\frac2x\cot(\ln x)~dx\] i spose since the lnx, that makes the right part simpler to deal with \[\ln x+2\ln(\sin(\ln x))\]

OpenStudy (psymon):

Okay, so I got it Im pretty sure, just forgot to keep the angle of sin as lnx.

OpenStudy (amistre64):

\[e^{\ln x+2\ln(\sin(\ln x))}w=C\] etc ... looks like a hoot

OpenStudy (amistre64):

you did fine

OpenStudy (amistre64):

since w=u', we can determine u; since we know y2 = u y1, we can then determine y2

OpenStudy (psymon):

Okay, cool. so....that means \[\mu'(x) = (wsin^{2}(lnx) + C)\]?

OpenStudy (psymon):

Or no, replace w with u', right

OpenStudy (amistre64):

solve for w

OpenStudy (amistre64):

yes, replace u' with the results for w

OpenStudy (psymon):

\[\mu' (x)\sin^{2}(lnx) = C\] \[\mu'(x) = C*\csc^{2}(lnx) \]

OpenStudy (psymon):

Integrate again, right?

OpenStudy (amistre64):

\[w=C\frac{\csc^2(\ln x)}{x}\] \[u'=C\frac{\csc^2(\ln x)}{x}\]

OpenStudy (amistre64):

integrate again, yes

OpenStudy (psymon):

Oops, x in there x_x

OpenStudy (psymon):

u = -C*cot(lnx) + C2

OpenStudy (amistre64):

looking good; c1 and c2 might clean it up more

OpenStudy (psymon):

Right. \[y _{2}=(-C _{1}\cot(lnx) + C _{2})xsin(lnx)\]

OpenStudy (psymon):

We need a C1 and C2, though, right?

OpenStudy (amistre64):

since c1 is arbitrary, then -c1 is redundant

OpenStudy (amistre64):

yes, y'' tells us that there are 2 arbitrary constants

OpenStudy (psymon):

Right, lol. I just didnt want to mess up, so I put it in still xD

OpenStudy (psymon):

Okay, so the book just decided to pick 2 numbers for C1 and C2. Does that mean they can be any 2 numbers?

OpenStudy (amistre64):

notice that y1 is incorporated into y2 at the moment, so the full solution would amount to y2

OpenStudy (amistre64):

yes; and if we had some initial conditiions we could then specify values of c1 and c2 as well

OpenStudy (psymon):

Yeah, I wasnt sure because usually textbooks would put something liek "answers may vary" in the backbut all of these have specific answers, so i wasnt sure.

OpenStudy (amistre64):

\[y = c_1 x \sin(\ln x)+c_2 x \cos(\ln x)\]

OpenStudy (psymon):

So I guess it doesnt matter what C1 and C2 I pick. Okay, cool. SO another thing I wasnt sure what to do with. Basically you do all the substituting and then replace u'' with w', but all I had left was \[w'e^{2x} = 0\]So how would I get my integrating factor out of that without a regular w term?

OpenStudy (amistre64):

theres no need to play with that .... our goal was in determining \(u\): \[u'=w\] \[u=\int w\]

OpenStudy (psymon):

Oh O.o Okay, cool, wouldnt have know. Alright, thanks a lot for taking thte time ^_^

OpenStudy (psymon):

\[xy'' + y' = 0; y _{1}=lnx\] \[y_{2} = \mu(x)lnx\] \[y_{2}' = \mu'(x)lnx + \frac{ \mu(x) }{ x }\] \[y''_{2}=\mu''(x)lnx + \frac{ \mu'(x) }{ x }+\frac{ x \mu'(x)-\mu(x) }{ x^{2} }\] \[x(\mu''(x)lnx + \frac{\mu'(x)}{x}+\frac{x \mu'(x)- \mu(x)}{x^{2}})+\mu'(x)lnx+ \frac{\mu(x)}{x}=0\] \[x \mu''(x)lnx + (2+lnx) \mu'(x)\] \[\mu''(x) = w'(x) = \frac{ dw }{ dx }\] \[\mu'(x) = w(x)\] \[(xlnx)\frac{dw}{dx}+(2+lnx)w = 0\] \[\frac{ dw }{ w }+\frac{ (2+lnx)dx }{ xlnx }=0\] \[\int\limits_{}^{}\frac{dw}{w}=-\int\limits_{}^{}\frac{ 2+lnx }{ xlnx }dx\] u = 2+lnx du = (1/x)dx dx = xdu \[\ln|w| = \int\limits_{}^{}\frac{ u }{ x(u-2) }*xdu\] \[\ln|w| = -\int\limits_{}^{} 1du+2\int\limits_{}^{}\frac{ 1 }{ u-2 }du\] \[\ln|w| =-u-2\ln|u-2|+C\] \[\ln|w| = -2-lnx-2\ln|lnx| + C\] Ill stop there. In the end, it says y2 is 1 in my book :/ So I want to make sure I did this lead-up stuff correct and then yeah, somehow figure out how the book gets 1, lol.

OpenStudy (amistre64):

\[xy'' + y' = 0;~y _{1}=lnx\] \[y=u~lnx\] \[y'=u'~lnx+ux^{-1}\] \[y''=u''~lnx+2u'x^{-1}-ux^{-2}\] --------------------- \[u''x~lnx+2u'-ux^{-1}+u'~lnx+ux^{-1}=0\] \[u''x~lnx+(2+lnx)u'=0\] ---------------------- \[w'x~lnx+(2+lnx)w=0\] \[w'+(\frac{2}{x~lnx}+\frac1x)w=0\] \[\frac 1w dw=-(\frac{2}{x~lnx}+\frac1x)dx \] \[ln~w=-2ln(ln~x)-ln~x+C \] \[w=C~\frac{1}{x~ln^{2}(x)} \] im up to here

OpenStudy (amistre64):

\[u = C_1\frac{1}{ln(x)}+C_2\]

OpenStudy (amistre64):

\[y=u lnx\] \[y=( C_1\frac{1}{ln(x)}+C_2) lnx\] \[y= C_1\frac{ln(x)}{ln(x)}+C_2 lnx\]